4. Thad Morgan, a motorcycle enthusiast, has been exploring the possibility of relaunching the Western Hombre brand of cycle that was popular in the 1930s. The retro-look cycle would be sold for $13,000 and at that price, Thad estimates 400 units would be sold each year. The variable cost to produce and sell the cycles would be $10,400 per unit. The annual fixed cost would be $988,000.

a. What is the break-even in unit sales?

b. What is the margin of safety in dollars? c. What is the degree of operating leverage? (Round your answer to 2 decimal places.)

5. Thad is worried about the selling price. Rumors are circulating that other retro brands of cycles may be revived. If so, the selling price for the Western Hombre would have to be reduced to $11,000 to compete effectively. In that event, Thad would also reduce fixed expenses to $751,000 by reducing advertising expenses, but he still hopes to sell 400 units per year.

d. What would the net operating income be in this situation? (Negative amount should be indicated by a minus sign.)
Answer #1

Answer

  • All working forms part of the answer
  • Requirement ‘a’

A

Sale price

$13,000

B

Variable cost

$10,400

C = A - B

Contribution margin per cycle

$2,600

D

Annual Fixed Cost

$988,000

E = D/C

Break even in unit sales

380

  • Requiremnet ‘b’

A

Estimated sales in units

400

B

Break even in unit sales

380

C = A - B

Margin of safety in units

20

D = C x $ 13000

Margin of safety in dollars

$260,000

  • Requirement ‘c’

A

Contribution margin per cycle

$2,600

B = 400 units x A

Total contribution margin

$1,040,000

C

Fixed Cost

$988,000

D = B - C

Net Operating Income

$52,000

E = B/D

Degree of Operating Leverage

20.00

  • Requirement ‘d’

A

Sale price

$11,000

B

Variable cost

$10,400

C = A - B

Contribution margin per cycle

$600

D = 400 units x C

Total contribution margin

$240,000

E

Fixed Cost

$751,000

F = D - E

Net Operating Income

($511,000)